OpenStudy (anonymous):

Estimate the sum or difference. 24 1/10 + 5 9/11 A 19 B 18 C 17

OpenStudy (studygurl14):

I think you meant to put -, not +

OpenStudy (studygurl14):

Because that gets 30, which isn't an answer choice

OpenStudy (studygurl14):

\(\large 24\frac{1}{10}+5\frac{9}{11}\rightarrow 24-6 =?\)

OpenStudy (anonymous):

I was thinking the same as @StudyGurl14 lol i been like is there option D lol!!

OpenStudy (anonymous):

its lois XD

OpenStudy (studygurl14):

lol yea @Cristel511